OpenStudy (anonymous):

find h(x) and g(x) such that f(x)=(h o g)(x) f(x)=[square root]8x+5 suppose that g(x)=8x+5

OpenStudy (anonymous):

im thinking that h(x) would be equal to \[\sqrt{x}\] but im not sure

OpenStudy (lgbasallote):

f(x) = h(g(x)) so... h(g(x)) = sqrt (8x + 5) g(x) = 8x + 5 therefore... \(\sqrt x\) is a logical answer

OpenStudy (lgbasallote):

you're right

OpenStudy (anonymous):

alright thank you

OpenStudy (lgbasallote):

welcome